# Advancing Brainwave Modeling with a Codebook-Based Foundation Model

### v1

@@ 5 Discussion
Foundation models have revolutionized numerous fields in computer science, enabling breakthroughs in natural language processing, computer vision and other domains. In this work, we introduce LaBraM++, an LBM with refined components compared to the original LaBraM model introduced in Jiang et al. (2024). By utilizing the benchmarking protocol described in Lee et al. (2025), our findings reveal that LaBraM++ significantly outperforms the original LaBraM architecture when trained on the same datasets, achieving a 6% improvement in overall performance. It also demonstrates a markedly lower training loss, suggesting more effective LBM optimization. When fully fine-tuned, LaBraM++ performs competitively against other pre-trained, open-source LBMs, many of which have been trained on substantially larger datasets.
In addition, the qualitative assessment of the reconstruction capabilities of LaBraM++’s tokenizer has shown that LaBraM++’s codebook effectively captures the overall trend and low-frequency components of the input signals with limitations in accurately reconstructing higher-frequency details. One promising avenue for enhancing LaBraM++’s performance lies in rethinking the signal reconstruction target. Instead of relying only on Fourier transform, which can miss short-term or localized frequency details, future versions of LaBraM++ could use scattering transform. Scattering transform preserves the time-frequency structure of signals, making it a suitable alternative for capturing richer signal characteristics. Additionally, incorporating training principles from the field of causal reasoning may further boost model performance. As highlighted in a recent work on causal modeling for LBMs Barmpas et al. (2024a), careful masking selection, both temporal and spatial, during the training phase is essential for capturing critical neural activity accurately during training.
@@ 6 Conclusion
In this work, we introduce LaBraM++, a Large Brainwave Foundation Model (LBM) built upon the original LaBraM introduced Jiang et al. (2024). The main contribution of this work is the design and training of an improved tokenizer, which in turn facilitates the development of a more effective LBM that shows improved performance across a range of tasks. Through a series of experiements, we showed that LaBraM++ significantly outperforms the original LaBraM architecture while achieving competitive performance against other pre-trained open-source LBMs.
In summary, LaBraM++ marks a step forward in the development of LBMs. Its improvements in performance and training efficiency, along with the potential for further enhancement through other signal representations and causal reasoning frameworks, set a strong foundation for future research in the area of LBMs.
@@ Appendix B Datasets
Emobrain Savran et al. [2006] BCI Competition IV-1 Blankertz et al. [2007] Self-Collected Dataset
SEED Zheng and Lu [2015] Grasp and Lift Luciw et al. [2014]
TUSL von Weltin et al. [2017] Inria BCI Challenge Margaux et al. [2012]
Self collected Jiang et al. [2024] Physionet MI Schalk et al. [2004]
Trujillo 2020 Trujillo [2020]
Trujillo 2017 Trujillo et al. [2017]
Siena Scalp Detti et al. [2020]
SPIS Resting Torkamani-Azar et al. [2019]
bi2015a Korczowski et al. [2019]
TUAR Buckwalter et al. [2021]
TUEP Veloso et al. [2017]
TUSZ Shah et al. [2018]

### v2

@@ 5 Discussion
In this work, we introduce LaBraM++, an LBM with refined components compared to the original LaBraM model introduced in Jiang et al., (2024). By utilizing the benchmarking protocol described in Lee et al., (2025), our findings reveal that LaBraM++ significantly outperforms the original LaBraM architecture when trained on the same datasets, achieving a 6% improvement in overall performance. When fully fine-tuned, LaBraM++ performs competitively against other pre-trained, open-source LBMs.
In addition, the qualitative assessment of the reconstruction capabilities of LaBraM++’s tokenizer has shown that LaBraM++’s codebook effectively captures the overall trend and low-frequency components of the input signals with limitations in accurately reconstructing higher-frequency details. One promising avenue for enhancing LaBraM++’s performance lies in rethinking the signal reconstruction target. Instead of relying only on Fourier transform, which can miss short-term or localized frequency details, future versions of LaBraM++ could use other additional transformations like scattering transform. Additionally, incorporating training principles from the field of causal reasoning may further boost model performance. As highlighted in a recent work on causal modeling for LBMs Barmpas et al., 2024a , careful masking selection, both temporal and spatial, during the training phase is essential for capturing critical neural activity accurately during training.
@@ 6 Conclusion
In this work, we introduce LaBraM++, a Large Brainwave Foundation Model (LBM) built upon the original LaBraM introduced Jiang et al., (2024). The main contribution of this work is the design and training of an improved tokenizer, which in turn facilitates the development of a more effective LBM that shows improved performance across a range of tasks. Through a series of experiements, we showed that LaBraM++ significantly outperforms the original LaBraM architecture while achieving competitive performance against other pre-trained open-source LBMs. In summary, LaBraM++ marks a step forward in the development of LBMs. Its improvements in performance and training efficiency, along with the potential for further enhancement through other signal representations and causal reasoning frameworks, set a strong foundation for future research in the area of LBMs.
@@ Appendix B Datasets
Emobrain Savran et al., (2006) BCI Competition IV-1 Blankertz et al., (2007) Self-Collected Dataset
SEED Zheng & Lu, (2015) Grasp and Lift Luciw et al., (2014)
TUSL von Weltin et al., (2017) Inria BCI Challenge Margaux et al., (2012)
Self collected Jiang et al., (2024) Physionet MI Schalk et al., (2004)
Trujillo 2020 Trujillo, (2020)
Trujillo 2017 Trujillo et al., (2017)
Siena Scalp Detti et al., (2020)
SPIS Resting Torkamani-Azar et al., (2019)
bi2015a Korczowski et al., (2019)
TUAR Buckwalter et al., (2021)
TUEP Veloso et al., (2017)
TUSZ Shah et al., (2018)